[Ncep.nhc.nco_contacts] TAFB Shift Log -- Technical Edition -- Tue Jul 25, 2017

TAFB Shift Log nhc.tsbadmin at noaa.gov
Wed Jul 26 02:30:00 UTC 2017


TAFB Shift Log for Tue Jul 25 2017 22:29:54 GMT-0400 (EDT)

Shift: Swing Shift

Pacific Forecaster: C McElroy

Atlantic Forecaster: G Rubio

Surface Analyst: M Formosa

Technical Issues Summary:

GFE/NCP issues:

N-AWIPS issues: 1500 L Shiner workstation locked up. 1530 L Matt Sardi  
investigating Shiner outage 1535 L EPAC 12Z NWPS run started 1419Z,  
finished 1719Z but no 12Z NWPS grids. Matt Sardi/Chris Mello investigating.  
1540 L Shiner back. Matt found loose mouse connection.

Other issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170726/a7cf0476/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list